# SPECIAL RELATIVITY

### v11

@@ 19.3.2 Why Did the Error in Insertion Device Theory Remain Undetected so Long?
We analyze the radiation associated with the first-order term in the expansion of Eq. (147) in powers of the small parameter $[K\theta_{x}\omega/(\gamma k_{w}c)]$ (see Eq. (150)). However, this approximation neglects the term $K\theta_{x}\omega\cos(k_{w}z^{\prime})/(\gamma k_{w}c)$, and as a result, all information about the transverse electron trajectory in the phase factor of Eq. (53) is lost. Under this approximation, the scale of the electron’s orbit is much smaller than the radiation’s diffraction size, and Eq. (60) yields fields that closely match those predicted by dipole radiation theory. Therefore, we consider the non-covariant approach sufficient for describing transverse electron motion in this context.
@@ 19.3.3 Influence of the Kick According to Conventional Theory
Therefore, introducing a deflection angle effectively corresponds to a rigid rotation of the entire system. When performing this rotation, it is important to note that the phase factor in Eq. (61)—which represents a spherical wavefront emanating from $z=0$—remains invariant under rotation. In contrast, the argument of the $\mathrm{sinc}(\cdot)$ function in Eq. (61) is altered by the rotation, as the point $(z_{0},0,0)$ is mapped to $(z_{0},-\eta_{x}z_{0},-\eta_{y}z_{0})$. As a result, after the rotation, Eq. (61) transforms accordingly:
@@ 19.3.4 Influence of the Kick According to Correct Coupling of Fields and Particles
To demonstrate this effect, we consider a covariant treatment that explicitly employs Lorentz transformations. Within this framework, introducing a kick leads to a non-zero red shift in the resonance frequency. This shift arises because, in Lorentz coordinates, the electron’s longitudinal velocity is reduced—from $v$ to $v-v\theta_{k}^{2}/2$while the speed of light remains constant at its electrodynamical value, $c$.
@@ 19.3.5 Experimental Test of SR Theory in 3rd Generation Light Source
One way to highlight the incompatibility between the standard approach to relativistic electrodynamics, which typically involves Maxwell’s equations, and the particle trajectories derived from non-covariant particle tracking is through a direct laboratory test of synchrotron radiation theory. Essentially, we are asserting that, despite extensive measurements accumulated over several decades, synchrotron radiation theory has yet to be definitively confirmed by experimental evidence.
@@ 19.4.1 Conventional Theory
[…]
@@
[…]
×$\times\exp\left\{{i\omega\left[{z^{\prime}\over{2\gamma^{2}c}}\left(1+\gamma^{2}\theta_{y}^{2}\right)+{z^{\prime 3}\over{6R^{2}c}}\right]}\right\}~{},$ (83)
@@ 19.4.3 Influence of the Kick According to Conventional Theory
[…]
@@
×$\times\left({1\over{\gamma^{2}}}+(\bar{\theta}_{y}-\eta_{y})^{2}+{(\bar{\theta}_{x}-\eta_{x})^{2}\over{3}}\right)$ (93)
@@
[…]
×$\times\exp\left\{{i\omega\left[{z^{\prime}\over{2\gamma^{2}c}}\left(1+\gamma^{2}(\bar{\theta}_{y}-\eta_{y})^{2}\right)+{z^{\prime 3}\over{6R^{2}c}}\right]}\right\}~{},$ (97)
@@
[…]
×$\times\exp\left\{{i\omega\left[{z^{\prime}\over{2\gamma^{2}c}}\left(1+\gamma^{2}\left(\theta_{y}-\eta_{y}\right)^{2}\right)+{z^{\prime 3}\over{6R^{2}c}}\right]}\right\}~{},$ (101)
@@ 19.5 Problem-Solving: Multiple Trajectory Kicks
It is important to note that, in principle, the entire history of the electron - from $t^{\prime}=-\infty$ to $t^{\prime}=\infty$must be considered, since the integral in Eq. (LABEL:revwied) extends over this range. However, the physical interpretation of this statement depends on the specific scenario. In practice, the integration should be restricted to the time intervals during which the electron contributes significantly to the radiation field.
@@
Although the integration in Eq.(43) is performed from $\infty$- ∞ to $\infty$, the only (edge) part of the trajectory into the bending magnets contributing to the integral is of the order of the radiation formation length $L_{f}$.
@@ 19.6.1 Existing Theory
[…]
@@ 20.1 Introductory Remarks
In the conventional framework, this wavefront tilt is treated as a physical phenomenon. That is, a transverse kick does not alter the orientation of the modulation wavefront, thereby reducing the efficiency of radiation in the direction of the beam motion  [^66] In standard XFEL operation, the electron beam is guided to remain closely aligned with the undulator axis. Nonetheless, random perturbations in the focusing system can introduce angular deviations or ”kicks.” The resultant misalignment between beam trajectory and wavefront normal has been discussed in the literature. It is traditionally understood that coherent radiation is emitted along the wavefront normal. Consequently, under the conventional field-particle coupling (which we argue is flawed), the mismatch between these directions diminishes radiation efficiency [79]. Studies of trajectory errors in XFEL amplification have shown that undulator magnetic fields must meet stringent tolerances. Interestingly, these tolerances have proven to be more stringent than necessary according to the corrected, four-dimensional (4D) covariant XFEL theory—an insight that helps explain the remarkable success of XFEL technology in recent decades..
@@ 20.2 Modulation Wavefront Orientation
We now examine the case where the electron beam is accelerated in the lab frame to acquire a small transverse velocity component $v_{y}$ along the $y$-axis. This raises an important question: how should synchronization be defined in the lab frame after this acceleration? Prior to the kicker setup, we chose a Lorentz coordinate system for the lab frame  [^67] Conventional XFEL theory adopts the absolute time convention to describe particle dynamics. Notably, employing the standard coupling of Maxwell’s equations with the corrected Newtonian equations to compute radiation from an ultrarelativistic electron beam in an undulator does not inherently introduce inaccuracies. In the case of straight-line motion—whether for the electron beam or the emitted undulator radiation—both covariant and non-covariant approaches produce identical trajectories. This ensures internal consistency between Maxwell’s equations and conventional particle tracking methods. However, this consistency brings to the forefront a critical question: how should synchronization be defined in the laboratory frame for a modulated electron beam upstream of the kicker setup? This is particularly relevant considering that the absolute time coordinatization was originally adopted in alignment with FEL theory. The most straightforward synchronization method employs the same set of synchronized clocks used in conventional particle tracking. This approach defines simultaneity based on light signals emitted from a stationary dipole source, assuming isotropic light propagation at speed $c$ in all directions.. After the beam acquires a small transverse velocity $v_{y}$, maintaining the original synchronization leads to complications in the electrodynamics of moving charges. As a result of such a boost, the transformation of time and spatial coordinates has the form of a Galilean transformation.
@@ 20.3 XFEL Radiation Configuration
where $k_{w}=2\pi/\lambda_{w}$ is the undulator wavenumber and $\vec{e}_{x,y}$ are unit vectors directed along the $x$ and $y$ axes. We neglected the transverse variation of the magnetic field. It is necessary to mention that in XFEL engineering we deal with very high-quality undulator systems, which have a sufficiently wide good-field-region, so that our studies, which refer to a simple model of undulator field nevertheless yield a correct quantitative description in a large variety of practical problems. The Lorentz force $\vec{F}=-e\vec{v}\times\vec{B}_{w}/c$ is used to derive the equation of motion of electrons with charge $-e$ and mass $m$ in the presence of magnetic field
@@
×(×$\vec{\nabla}\times(\vec{\nabla}\times{\vec{E}})=\vec{\nabla}(\vec{\nabla}\cdot{\vec{E}})-\nabla^{2}\vec{E}$ (118)
@@
Here follows an explanation of the resonance condition which is elementary in the sense that we can see what is happening physically. The field of electromagnetic wave has only transverse components, so the energy exchange between the electron and electromagnetic wave is due to a transverse component of the electron velocity. For effective energy exchange between the electron and the wave, the scalar product $-e\vec{v}_{\perp}\cdot\vec{E}$ should be kept nearly constant along the whole undulator length. We see that required synchronism $k_{w}+\omega/c-\omega/v_{z}=0$ takes place when the wave advances the electron beam by the wavelength at one undulator period $\lambda_{w}/v_{z}=\lambda/(c-v_{z})$, where $\lambda=2\pi/\omega$ is the radiation wavelength. This tells us that the angle between the transverse velocity of the particle $\vec{v}_{\perp}$ and the vector of the electric field $\vec{E}$ remains nearly constant. Since $v_{z}\simeq c$ this resonance condition may be written as $\lambda\simeq\lambda_{w}/(2\gamma_{z}^{2})=\lambda_{w}(1+K^{2})/(2\gamma^{2})$.
@@
[…]
@@
To understand the angular distribution, consider a triangle formed by a radiation path at a small angle $\theta$, with altitude $r\simeq\theta z$ and base $z$. The diagonal path length $s$ exceeds the base by an amount $\Delta=s-z\simeq z\theta^{2}/2$. When $\Delta$ equals one radiation wavelength, destructive interference occurs, as the phase contributions from different oscillators become uniformly distributed over 00 to $2\pi$.
@@ 20.6 Experimental Test: Electrodynamics - Dynamics Coupling
20.7 Debunking the Modulation Wavefront Tilt Myth
In conventional XFEL theory, wavefront tilt is typically regarded as a physically real effect. However, a common misconception in accelerator physics relates to the interpretation of this tilt. In the ultrarelativistic regime, wavefront tilt does not have an exact, objective meaning. Its value depends on the choice of clock synchronization procedure in the lab frame and, consequently, can be assigned arbitrarily within the interval $(0,\theta_{k})$.
@@ 21 Appendix
We start with the solution of Maxwell’s equation in the space-time domain, the well-known Lienard-Wiechert expression, and we subsequently apply a Fourier transformation. The Lienard-Wiechert expression for the electric field of a point charge $(-e)$ reads (see, e.g. [41]):
@@
[…]
@@
$\vec{\bar{E}}$ (((( $\vec{r}_{o},\omega)=-{e}\int_{-\infty}^{\infty}dt^{\prime}{\vec{n}\over{|\vec{r}_{o}-\vec{r^{\prime}}(t^{\prime})|^{2}}}\exp\left[i\omega\left(t^{\prime}+{|\vec{r}_{o}-\vec{r^{\prime}}(t^{\prime})|\over{c}}\right)\right]$ (130)
[…]
@@
[…]
×$\times\exp\left\{i\omega\left(t^{\prime}+{|\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime})|\over{c}}\right)\right\}~{}.$
@@
[…]
×$\times\exp\left\{i\omega\left(t^{\prime}+{|\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime})|\over{c}}\right)\right\}~{}.$
@@
×$\times\int_{-L/2}^{L/2}dz^{\prime}\left\{\frac{K}{2i\gamma}\left[\exp\left(2ik_{w}z^{\prime}\right)-1\right]\vec{e}_{x}+\vec{\theta}\exp\left(ik_{w}z^{\prime}\right)\right\}$ (139)
×$\times\exp\left[i\left(C+{\omega\theta^{2}\over{2c}}\right)z^{\prime}-{K\theta_{x}\over{\gamma}}{\omega\over{k_{w}c}}\cos(k_{w}z^{\prime})-{K^{2}\over{8\gamma^{2}}}{\omega\over{k_{w}c}}\sin(2k_{w}z^{\prime})\right]~{}.$ (140)
@@
×$\times\int_{-L/2}^{L/2}dz^{\prime}\exp\left[i\left(C+{\omega\theta^{2}\over{2c}}\right)z^{\prime}\right]$ (145)
×$\times\left\{\frac{K}{2i\gamma}\left[\exp\left(2ik_{w}z^{\prime}\right)-1\right]\vec{e}_{x}+\vec{\theta}\exp\left(ik_{w}z^{\prime}\right)\right\}\exp\left[i(n+2m)k_{w}z^{\prime}\right]~{},$ (146)
@@
[…]
@@
Here we are dealing specifically with the first harmonic. Therefore, these extra-simplifications apply. We neglect both the term in $\cos(k_{w}z^{\prime})$ in the phase of Eq. (141) and the term in $\vec{\theta}$ in Eq. (141). First, non-negligible terms in the expansion of $J_{n}(v)$ are those for small values of $n$, since $J_{n}(v)\sim v^{n}$, with $|v|\ll 1$. The value $n=0$ gives a non-negligible contribution $J_{0}(v)\sim 1$. Then, since the integration in $dz^{\prime}$ is performed over a large number of undulator periods $N_{w}\gg 1$, all terms of the expansion in Eq. (147) but those for $m=-1$ and $m=0$ average to zero due to resonance approximation. Note that surviving contributions are proportional to $K/\gamma$, and can be traced back to the term in $\vec{e}_{x}$ only, while the term in $\vec{\theta}$ in Eq. (147) averages to zero for $n=0$. Values $n=\pm 1$ already give negligible contributions. In fact, $J_{\pm 1}(v)\sim v$. Then, the term in $\vec{e}_{x}$ in Eq. (147) is $v$ times the term with $n=0$ and is immediately negligible, regardless of the values of $m$. The term in $\vec{\theta}$ would survive averaging when $n=1,~{}m=-1$ and when $n=-1,~{}m=0$. However, it scales as $\vec{\theta}v$. Now, using condition $|\Delta\omega|/\omega_{r}\ll 1$ we see that, for observation angles of interest $\theta^{2}\ll 1/\gamma_{z}^{2}$, $|\vec{\theta}|~{}|v|\sim(\sqrt{2}~{}K~{}/\sqrt{2+K^{2}}~{})~{}\bar{\gamma}_{z}\theta^{2}\ll K/\gamma$. Therefore, the term in $\vec{\theta}$ is negligible with respect to the term in $\vec{e}_{x}$ for $n=0$, that scales as $K/\gamma$. All terms corresponding to larger values of $|n|$ are negligible.
Summing up, all terms of the expansion in Eq. (143) but those for $n=0$ and $m=-1$ or $m=0$ give negligible contribution. After definition of
@@
[…]
@@
[…]
@@
[…]

### v12

@@ 20.3.2 Why Did the Error in Insertion Device Theory Remain Undetected so Long?
We analyze the radiation associated with the first-order term in the expansion of Eq. (147) in powers of the small parameter $K\theta_{x}\omega/(\gamma k_{w}c)]$ (see Eq. (150)). However, this approximation neglects the term $K\theta_{x}\omega\cos(k_{w}z^{\prime})/(\gamma k_{w}c)$, and as a result, all information about the transverse electron trajectory in the phase factor of Eq. (53) is lost. Under this approximation, the scale of the electron’s orbit is much smaller than the radiation’s diffraction size, and Eq. (60) yields fields that closely match those predicted by dipole radiation theory. Therefore, we consider the non-covariant approach sufficient for describing transverse electron motion in this context.
@@ 20.3.3 Influence of the Kick According to Conventional Theory
Therefore, introducing a deflection angle effectively corresponds to a rigid rotation of the entire system. When performing this rotation, it is important to note that the phase factor in Eq. (61)—which represents a spherical wavefront emanating from z=0z=0—remains invariant under rotation. In contrast, the argument of the $\mathrm{sinc}(\cdot)$ function in Eq. (61) is altered by the rotation, as the point $(z_{0},0,0)$ is mapped to $(z_{0},-\eta_{x}z_{0},-\eta_{y}z_{0})$. As a result, after the rotation, Eq. (61) transforms accordingly:
@@ 20.3.4 Influence of the Kick According to Correct Coupling of Fields and Particles
To demonstrate this effect, we consider a covariant treatment that explicitly employs Lorentz transformations. Within this framework, introducing a kick leads to a non-zero red shift in the resonance frequency. This shift arises because, in Lorentz coordinates, the electron’s longitudinal velocity is reduced—from $v$ to $v-v\theta_{k}^{2}/2$ -while the speed of light remains constant at its electrodynamical value, $c$.
@@ 20.3.5 Experimental Test of SR Theory in 3rd Generation Light Source
One way to highlight the incompatibility between the standard approach to relativistic electrodynamics, which typically involves Maxwell’s equations, and the particle trajectories derived from non-covariant particle tracking is through a direct laboratory test of synchrotron radiation theory.
@@ 20.4.1 Conventional Theory
$\vec{\widetilde{E}}={i\omega e\over{c^{2}z_{0}}}\int_{-\infty}^{\infty}dz^{\prime}{e^{i\Phi_{T}}}\left({z^{\prime}+R\theta_{x}\over{R}}\vec{e_{x}}+\theta_{y}\vec{e_{y}}\right)$~ (77)
@@
$\vec{\widetilde{E}}={i\omega e\over{c^{2}z_{0}}}e^{i\Phi_{s}}e^{i\Phi_{0}}\int_{-\infty}^{\infty}dz^{\prime}\left({z^{\prime}\over{R}}\vec{e_{x}}+\theta_{y}\vec{e_{y}}\right)$ (82)
$\times\exp\left\{{i\omega\left[{z^{\prime}\over{2\gamma^{2}c}}\left(1+\gamma^{2}\theta_{y}^{2}\right)+{z^{\prime 3}\over{6R^{2}c}}\right]}\right\}~,$ (83)
@@ 20.4.3 Influence of the Kick According to Conventional Theory
$\vec{\widetilde{E}}={i\omega e\over{c^{2}z_{0}}}\int_{-\infty}^{\infty}dz^{\prime}{e^{i\Phi_{T}}}\left({z^{\prime}+R(\bar{\theta}_{x}-\eta_{x})\over{R}}\vec{e_{x}}+{(\bar{\theta}_{y}-\eta_{y})}\vec{e_{y}}\right)$~ (89)
@@
$\times\left({1\over{\gamma^{2}}}+(\bar{\theta}_{y}-\eta_{y})^{2}+{(\bar{\theta}_{x}-\eta_{x})^{2}\over{3}}\right)$ (93)
@@
$\vec{\widetilde{E}}={i\omega e\over{c^{2}z_{0}}}e^{i\Phi_{s}}e^{i\Phi_{0}}\int_{-\infty}^{\infty}dz^{\prime}\left({z^{\prime}\over{R}}\vec{e_{x}}+(\bar{\theta}_{y}-\eta_{y})\vec{e_{y}}\right)$ (96)
$\times\exp\left\{{i\omega\left[{z^{\prime}\over{2\gamma^{2}c}}\left(1+\gamma^{2}(\bar{\theta}_{y}-\eta_{y})^{2}\right)+{z^{\prime 3}\over{6R^{2}c}}\right]}\right\}~,$ (97)
@@
$\vec{\widetilde{E}}={i\omega e\over{c^{2}z_{0}}}e^{i\Phi_{s}}e^{i\Phi_{0}}\int_{-\infty}^{\infty}dz^{\prime}\left({z^{\prime}\over{R}}\vec{e_{x}}+\left(\theta_{y}-\eta_{y}\right)\vec{e_{y}}\right)$ (100)
$\times\exp\left\{{i\omega\left[{z^{\prime}\over{2\gamma^{2}c}}\left(1+\gamma^{2}\left(\theta_{y}-\eta_{y}\right)^{2}\right)+{z^{\prime 3}\over{6R^{2}c}}\right]}\right\}~,$ (101)
@@ 20.5 Problem-Solving: Multiple Trajectory Kicks
It is important to note that, in principle, the entire history of the electron - from $t^{\prime}=$-$\infty$ to $t^{\prime}=\infty$ -must be considered, since the integral in Eq. (LABEL:revwied) extends over this range. However, the physical interpretation of this statement depends on the specific scenario. In practice, the integration should be restricted to the time intervals during which the electron contributes significantly to the radiation field.
@@
Although the integration in Eq.(43) is performed from -$\infty$ to $\infty$, the only (edge) part of the trajectory into the bending magnets contributing to the integral is of the order of the radiation formation length $L_{f}$.
@@ 20.6.1 Existing Theory
$\int_{0}^{\infty}x\sin[(3/2)\alpha(x+x^{3}/3)]dx=(1/\sqrt{3})K_{2/3}(\alpha)~,$ (106)
$\int_{0}^{\infty}\cos[(3/2)\alpha(x+x^{3}/3)]dx=(1/\sqrt{3})K_{1/3}(\alpha)~.$ (107)
@@ 21.1 Introductory Remarks
In the conventional framework, this wavefront tilt is treated as a physical phenomenon. That is, a transverse kick does not alter the orientation of the modulation wavefront, thereby reducing the efficiency of radiation in the direction of the beam motion.  [^68] In standard XFEL operation, the electron beam is guided to remain closely aligned with the undulator axis. Nonetheless, random perturbations in the focusing system can introduce angular deviations or ”kicks.” The resultant misalignment between beam trajectory and wavefront normal has been discussed in the literature. It is traditionally understood that coherent radiation is emitted along the wavefront normal. Consequently, under the conventional field-particle coupling (which we argue is flawed), the mismatch between these directions diminishes radiation efficiency [80]. Studies of trajectory errors in XFEL amplification have shown that undulator magnetic fields must meet stringent tolerances. Interestingly, these tolerances have proven to be more stringent than necessary according to the corrected, four-dimensional (4D) covariant XFEL theory—an insight that helps explain the remarkable success of XFEL technology in recent decades.
@@ 21.2 Modulation Wavefront Orientation
We now examine the case where the electron beam is accelerated in the lab frame to acquire a small transverse velocity component $v_{y}$ along the $y$-axis. This raises an important question: how should synchronization be defined in the lab frame after this acceleration? Prior to the kicker setup, we chose a Lorentz coordinate system for the lab frame.  [^69] Conventional XFEL theory adopts the absolute time convention to describe particle dynamics. Notably, employing the standard coupling of Maxwell’s equations with the corrected Newtonian equations to compute radiation from an ultrarelativistic electron beam in an undulator does not inherently introduce inaccuracies. In the case of straight-line motion—whether for the electron beam or the emitted undulator radiation—both covariant and non-covariant approaches produce identical trajectories. This ensures internal consistency between Maxwell’s equations and conventional particle tracking methods. However, this consistency brings to the forefront a critical question: how should synchronization be defined in the laboratory frame for a modulated electron beam upstream of the kicker setup? This is particularly relevant considering that the absolute time coordinatization was originally adopted in alignment with FEL theory. The most straightforward synchronization method employs the same set of synchronized clocks used in conventional particle tracking. This approach defines simultaneity based on light signals emitted from a stationary dipole source, assuming isotropic light propagation at speed $c$ in all directions. After the beam acquires a small transverse velocity $v_{y}$, maintaining the original synchronization leads to complications in the electrodynamics of moving charges. As a result of such a boost, the transformation of time and spatial coordinates has the form of a Galilean transformation.
@@ 21.3 XFEL Radiation Configuration
where $k_{w}=2\pi/\lambda_{w}$ is the undulator wavenumber and $\vec{e}_{x,y}$ are unit vectors directed along the $x$ and $y$ axes. We neglected the transverse variation of the magnetic field. It is necessary to mention that in XFEL engineering we deal with very high-quality undulator systems, which have a sufficiently wide good-field-region, so that our studies, which refer to a simple model of undulator field nevertheless yield a correct quantitative description in a large variety of practical problems. The Lorentz force $\vec{F}=-e\vec{v}\times\vec{B}_{w}/c$ is used to derive the equation of motion of electrons with charge e-e and mass $m$ in the presence of magnetic field
@@
$\vec{\nabla}\times(\vec{\nabla}\times{\vec{E}})=\vec{\nabla}(\vec{\nabla}\cdot{\vec{E}})-\nabla^{2}\vec{E}$ (118)
@@
Here follows an explanation of the resonance condition which is elementary in the sense that we can see what is happening physically. The field of electromagnetic wave has only transverse components, so the energy exchange between the electron and electromagnetic wave is due to a transverse component of the electron velocity. For effective energy exchange between the electron and the wave, the scalar product $e\vec{v}_{\perp}\cdot\vec{E}$ should be kept nearly constant along the whole undulator length. We see that required synchronism $k_{w}+\omega/c-\omega/v_{z}=0$ takes place when the wave advances the electron beam by the wavelength at one undulator period $\lambda_{w}/v_{z}=\lambda/(c-v_{z})$, where $\lambda=2\pi/\omega$ is the radiation wavelength. This tells us that the angle between the transverse velocity of the particle $\vec{v}_{\perp}$ and the vector of the electric field $\vec{E}$ remains nearly constant. Since $v_{z}\simeq c$ this resonance condition may be written as $\lambda\simeq\lambda_{w}/(2\gamma_{z}^{2})=\lambda_{w}(1+K^{2})/(2\gamma^{2})$.
@@
$\vec{\bar{E}}=\int d\vec{r^{\prime}}~G(\vec{r},\vec{r^{\prime}})\left[-4\pi i\frac{\omega}{c^{2}}\vec{\bar{j}_{\perp}}\right]~.$ (124)
@@
To understand the angular distribution, consider a triangle formed by a radiation path at a small angle $\theta$, with altitude $r\simeq\theta z$ and base $z$. The diagonal path length $s$ exceeds the base by an amount $\Delta=s-z\simeq z\theta^{2}/2$. When $\Delta$ equals one radiation wavelength, destructive interference occurs, as the phase contributions from different oscillators become uniformly distributed over 0 to $2\pi$.
@@ 21.6 Experimental Test: Electrodynamics - Dynamics Coupling
21.7 Wavefront Tilt and Degradation of Electron Beam Modulation
In conventional XFEL theory, wavefront tilt is often treated as a physically meaningful and measurable quantity. However, a common misconception in accelerator physics concerns the interpretation of this tilt. In the ultrarelativistic regime, wavefront tilt lacks a unique, objective definition. Its value depends on the choice of clock synchronization convention in the laboratory frame and can therefore be assigned arbitrarily within the interval $(0,\theta_{k})$.
@@ 22 Appendix
We start with the solution of Maxwell’s equation in the space-time domain, the well-known Lienard-Wiechert expression, and we subsequently apply a Fourier transformation. The Lienard-Wiechert expression for the electric field of a point charge (e)(-e) reads (see, e.g. [41]):
@@
$\vec{\bar{E}}(\vec{r}_{o},\omega)$ $=$ $-e\int_{-\infty}^{\infty}dt^{\prime}{\vec{n}-\vec{\beta}\over{\gamma^{2}(1-\vec{n}\cdot\vec{\beta})^{2}|\vec{r}_{o}-\vec{r^{\prime}}|^{2}}}\exp\left[i\omega\left(t^{\prime}+{|\vec{r}_{o}-\vec{r^{\prime}}(t^{\prime})|\over{c}}\right)\right]$ (127)
$-{e\over{c}}\int_{-\infty}^{\infty}dt^{\prime}{\vec{n}\times[(\vec{n}-\vec{\beta})\times\dot{\vec{\beta}}]\over{(1-\vec{n}\cdot\vec{\beta})^{2}|\vec{r}_{o}-\vec{r^{\prime}}|}}\exp\left[i\omega\left(t^{\prime}+{|\vec{r}_{o}-\vec{r^{\prime}}(t^{\prime})|\over{c}}\right)\right].$
@@
$\vec{\bar{E}}$ $($ $\vec{r}_{o},\omega)=-{e}\int_{-\infty}^{\infty}dt^{\prime}{\vec{n}\over{|\vec{r}_{o}-\vec{r^{\prime}}(t^{\prime})|^{2}}}\exp\left[i\omega\left(t^{\prime}+{|\vec{r}_{o}-\vec{r^{\prime}}(t^{\prime})|\over{c}}\right)\right]$ (130)
$+{e\over{c}}\int_{-\infty}^{\infty}dt^{\prime}{d\over{dt^{\prime}}}\left[{\vec{\beta}-\vec{n}\over{({1-\vec{n}\cdot\vec{\beta}})|\vec{r}_{o}-\vec{r^{\prime}}(t^{\prime})|}}\right]\exp\left[i\omega\left(t^{\prime}+{|\vec{r}_{o}-\vec{r^{\prime}}(t^{\prime})|\over{c}}\right)\right].$
@@
$\vec{\bar{E}}(\vec{r_{o}},\omega)$ $=$ $-{i\omega e\over{c}}\int_{-\infty}^{\infty}dt^{\prime}\left[{\vec{\beta}-\vec{n}\over{|\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime})|}}-{ic\over{\omega}}{\vec{n}\over{|\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime})|^{2}}}\right]$ (133)
$\times\exp\left\{i\omega\left(t^{\prime}+{|\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime})|\over{c}}\right)\right\}~.$
@@
$\vec{\bar{E}}(\vec{r_{o}},\omega)$ $=$ $-{i\omega e\over{c}}\int_{-\infty}^{\infty}dt^{\prime}\left[-{\vec{n}\times(\vec{n}\times{\vec{\beta}})\over{|\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime})|}}+{i{c}\over{\omega}}{\vec{\beta}-\vec{n}-2\vec{n}(\vec{n}\cdot\vec{\beta})\over{|\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime})|^{2}}}\right]$ (137)
$\times\exp\left\{i\omega\left(t^{\prime}+{|\vec{r_{o}}-\vec{r^{\prime}}(t^{\prime})|\over{c}}\right)\right\}~.$
@@
$\times\int_{-L/2}^{L/2}dz^{\prime}\left\{\frac{K}{2i\gamma}\left[\exp\left(2ik_{w}z^{\prime}\right)-1\right]\vec{e}_{x}+\vec{\theta}\exp\left(ik_{w}z^{\prime}\right)\right\}$ (139)
$\times\exp\left[i\left(C+{\omega\theta^{2}\over{2c}}\right)z^{\prime}-{K\theta_{x}\over{\gamma}}{\omega\over{k_{w}c}}\cos(k_{w}z^{\prime})-{K^{2}\over{8\gamma^{2}}}{\omega\over{k_{w}c}}\sin(2k_{w}z^{\prime})\right]~.$ (140)
@@
$\times\int_{-L/2}^{L/2}dz^{\prime}\exp\left[i\left(C+{\omega\theta^{2}\over{2c}}\right)z^{\prime}\right]$ (145)
$\times\left\{\frac{K}{2i\gamma}\left[\exp\left(2ik_{w}z^{\prime}\right)-1\right]\vec{e}_{x}+\vec{\theta}\exp\left(ik_{w}z^{\prime}\right)\right\}\exp\left[i(n+2m)k_{w}z^{\prime}\right]~,$ (146)
@@
This means that, independently of $K$, $|v|\ll 1$ and we may expand $J_{n}(v)$ in Eq. (147) according to $J_{n}(v)\simeq[2^{-n}/\Gamma(1+n)]~v^{n}$, $\Gamma(\cdot)$ being the Euler gamma function $\Gamma(z)=\int_{0}^{\infty}dt~t^{z-1}\exp[-t]$
@@
Here we are dealing specifically with the first harmonic. Therefore, these extra-simplifications apply. We neglect both the term in $\cos(k_{w}z^{\prime})$ in the phase of Eq. (141) and the term in $\vec{\theta}$ in Eq. (141). First, non-negligible terms in the expansion of $J_{n}(v)$ are those for small values of $n$, since $J_{n}(v)\sim v^{n}$, with $|v|\ll 1$. The value $n=0$ gives a non-negligible contribution $J_{0}(v)\sim 1$. Then, since the integration in $dz^{\prime}$ is performed over a large number of undulator periods $N_{w}\gg 1$, all terms of the expansion in Eq. (147) but those for m=1m=-1 and $m=0$ average to zero due to resonance approximation. Note that surviving contributions are proportional to $K/\gamma$, and can be traced back to the term in $\vec{e}_{x}$ only, while the term in $\vec{\theta}$ in Eq. (147) averages to zero for $n=0$. Values $n=\pm 1$ already give negligible contributions. In fact, $J_{\pm 1}(v)\sim v$. Then, the term in $\vec{e}_{x}$ in Eq. (147) is $v$ times the term with $n=0$ and is immediately negligible, regardless of the values of $m$. The term in $\vec{\theta}$ would survive averaging when n=1,m=1n=1,~m=-1 and when n=1,m=0n=-1,~m=0. However, it scales as $\vec{\theta}v$. Now, using condition $|\Delta\omega|/\omega_{r}\ll 1$ we see that, for observation angles of interest $\theta^{2}\ll 1/\gamma_{z}^{2}$, $|\vec{\theta}|~|v|\sim(\sqrt{2}~K~/\sqrt{2+K^{2}}~)~\bar{\gamma}_{z}\theta^{2}\ll K/\gamma$. Therefore, the term in $\vec{\theta}$ is negligible with respect to the term in $\vec{e}_{x}$ for $n=0$, that scales as $K/\gamma$. All terms corresponding to larger values of $|n|$ are negligible.
Summing up, all terms of the expansion in Eq. (143) but those for $n=0$ and m=1m=-1 or $m=0$ give negligible contribution. After definition of
@@
${\vec{\widetilde{E}}}=-\frac{K\omega e}{2c^{2}z_{0}\gamma}A_{JJ}\exp\left[i\frac{\omega\theta^{2}z_{0}}{2c}\right]\int_{-L/2}^{L/2}dz^{\prime}\exp\left[i\left(C+{\omega\theta^{2}\over{2c}}\right)z^{\prime}\right]\vec{e}_{x}~.$ (152)
@@
$s(z)=\int_{0}^{z}dz^{\prime}\left[1+\left(\frac{dx}{dz^{\prime}}\right)^{2}+\left(\frac{dy}{dz^{\prime}}\right)^{2}\right]^{1/2}$ (156)
$\simeq\int_{0}^{z}dz^{\prime}\left[1+\frac{1}{2}\left(\frac{dx_{r}}{dz^{\prime}}\right)^{2}+\frac{1}{2}\left(\frac{dy_{r}}{dz^{\prime}}\right)^{2}+\frac{1}{2}\left(\eta_{x}^{2}+\eta_{y}^{2}\right)+\eta_{x}\frac{dx_{r}}{dz^{\prime}}+\eta_{y}\frac{dy_{r}}{dz^{\prime}}\right]$ (157)
@@
where we expanded the square root around unity in the first passage, we made use of Eq. (155), and of the fact that the curvilinear abscissa along the reference path is $s_{r}(z)\simeq z+\int_{0}^{z}dz^{\prime}[(dx_{r}/dz^{\prime})^{2}/2+(dy_{r}/dz^{\prime})^{2}/2]$.